how do u draw a rectangle with the perimeter of 12 cm and the area of 9 sq cm?
Answer by mananth(16946)

2(l+w)= perimeter
2(l+w)=12
/2
l+w = 6
Area = l*w
l*w = 9
w= 9/l
..
substitute w
l+w =6
l+9/l = 6
l^2+9=6l
l^2-6l+9=0
(l-3)^2=0
l=3
w=3
Its a square